OpenStudy (anonymous):

What is the 17th term of the arithmetic sequence 17, 21, 25, … ?

OpenStudy (ash2326):

nth term of an arithmetic sequence \(a_n\) is given as \[a_n=a_1+(n-1)d\] \[a_1 =\text{first term of the sequence}\] \[d= \text{common difference}\] What is \(a_1\) and \(d\) here?

OpenStudy (anonymous):

a1= 17 d=4

OpenStudy (ash2326):

We need to find the 17th term, so what is \(a_n\) here? and what's n?

OpenStudy (anonymous):

an is 17

OpenStudy (ash2326):

Let's substitute these \[a_{17}=a_1+(n-1)d\] \[a_{17}=17+(17-1) \times 4\] Can you solve this?

OpenStudy (anonymous):

81

OpenStudy (ash2326):

good:D Did you understand how we solved this?

OpenStudy (anonymous):

yeah, kinda thank youu

OpenStudy (ash2326):

@CCarmichael18 \[\large\mathcal{Welcome\ to\ Open\ Study\ :)}\] If you ever need help here, do let me know:)